10. A firm supplies 200 units of a good at a price of Rs 5 per unit. When

price changes it supplies 100 units less. Price Elasticity of Supply is 2.5.

Calculate price after change in quantity.


1
Expert's answer
2021-03-08T09:23:35-0500

"Q=200" ",P=200\times 5=Rs.1000"

"\Delta Q=100-200=-100"

let price after change in quantity"=x"

"\Delta P=x-1000"

"E_s=\Delta Q/ \Delta P\times P/Q"

"2.5=(-100)/(x-1000)\times 1000/200"

"2.5(x-1000)=-500"

"2.5x-1000=-500"

"2.5x=500"

"x=Rs200"

price per unit "=200\/100=Rs 2"